OpenStudy (anonymous):

(2xy)(4y) =2*4=8 = x*x*y=x^2y 8x^2y Did I simplify this right?

hartnn (hartnn):

2xy * 4y 2*4 = 8 is correct but we have 2 "y"s its x*y*y = xy^2 (and not x^2y)

OpenStudy (anonymous):

Okay, thanks.